OpenStudy (anonymous):

Please Help! Will fan and medal Find an equation for the line that passes through (–3, 8) and perpendicular to y =-1/4x -3

OpenStudy (jdoe0001):

hmmm what's the slope of -> =-1/4x -3 you think?

OpenStudy (anonymous):

-1/4

OpenStudy (jdoe0001):

so the slope of a line perpendicular to that one, will have a NEGATIVE RECIPROCAL slope so this one is \(\bf -\cfrac{1}{{\color{blue}{ 4}}}\qquad negative\to +\cfrac{1}{{\color{blue}{ 4}}}\qquad reciprocal\to +\cfrac{{\color{blue}{ 4}}}{1}\to 4\) so you're really being asked to find the equation of a line with a slope of "4" and that passes through (-3, 8) thus \(\bf \begin{array}{lllll} &x_1&y_1\ &({\color{red}{ -3}}\quad ,&{\color{blue}{ 8}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= 4 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (jdoe0001):

hmm anyhow \(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ -3}}\quad ,&{\color{blue}{ 8}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= 4 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values }\\ \qquad \uparrow\\ \textit{point-slope form}\) well, you're not asked to set it in slope-intercept form.... so you can just leave it like that in point-slope anyhow

OpenStudy (anonymous):

Thanks :D

OpenStudy (jdoe0001):

yw